OpenStudy (anonymous):

in shadow bay in july, high tide is at 1:30pm. the water level is 4 feet at high tide and 0 feet at low tide. Assuming the next high tide is eacactly 12 hours later and the height of the water can be modeled by a cosine curve, find an equation for shadow bay's water level in july as a function of time (t)

OpenStudy (anonymous):

@ribhu

OpenStudy (anonymous):

Do you get it?

OpenStudy (ribhu):

yeah kind of i am trying.

OpenStudy (anonymous):

I'll post my answer choices

OpenStudy (anonymous):

F(t)=2cospi/6t+1 F(t)=2cospi/6t+2 F(t)=4cospi/2t+1 F(t)=4cospi/2+2

OpenStudy (anonymous):

?

OpenStudy (ribhu):

b

OpenStudy (anonymous):

How?

OpenStudy (ribhu):

let at 1:30 pm t=0

OpenStudy (ribhu):

height would be 4

OpenStudy (ribhu):

at 1:30 am t= 12hrs, height = 4

OpenStudy (anonymous):

Ohhhh never mond

OpenStudy (anonymous):

*mind lol

OpenStudy (anonymous):

Thanks pal!